OpenStudy (anonymous):

okay, soo, i have a equation here that i can't seem to get the right answer for, please help ! (square root of 2x+5)-(square root of x-1)=(square root of x+2). i don't know how to make the square root sign on my keyboard sorry, please help !

OpenStudy (anonymous):

\[\sqrt{2x+5}-\sqrt{x-1}=\sqrt{x+2}\]

OpenStudy (anonymous):

\[\sqrt{2x+5}-\sqrt{x-1}=\sqrt{x+2}\]

OpenStudy (anonymous):

yes there is the formula, please help me!

OpenStudy (anonymous):

square both sides but don't distribute what do you get?

OpenStudy (anonymous):

\[(\sqrt{2x+5}-\sqrt{x-1})^2=x+2\]?

OpenStudy (anonymous):

\[(\sqrt{2x+5}-\sqrt{x-1})^{2}=x\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

x+2

OpenStudy (anonymous):

but yes

OpenStudy (anonymous):

x+2 sorry, miss type

OpenStudy (anonymous):

now lets write the square root in power form, i think that will help

OpenStudy (anonymous):

\[(2x+5)^{1/2}-(x-1)^{1/2})^2=x+2\]

OpenStudy (anonymous):

hmm, okay i see what you did there

OpenStudy (anonymous):

ehh that looks a little more worst lets just go back and workt the foil out with square roots \[(\sqrt{2x+5}-\sqrt{x-1})(\sqrt{2x+5}-\sqrt{x-1})\]

OpenStudy (anonymous):

we know \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

yes, squaring a binomial, the 3 step rule

OpenStudy (anonymous):

square the first term, multiply the first and second to get the second term then double it, then square the last term

OpenStudy (anonymous):

foiling you get \[2x+5-2(\sqrt{2x+5})(\sqrt{x-1})+x-1=x+2\]

OpenStudy (anonymous):

i think i can answer this question, i just need you to verify a few things with me first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for an example when you have \[(2\sqrt{2x+5}-x+2)(2\sqrt{2x+5-x+2})\] and you are simplfilying and multiplying it through, would the first part of it be \[4\sqrt{4x ^{2}+20x+25}\] ?

OpenStudy (anonymous):

oops, the second one isn't supposed to have a radical over everything, just the 2x+5 like the first one

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that would be how to multiply